# A Survey on Large Language Models from General Purpose to Medical Applications: Datasets, Methodologies, and Evaluations

### v1

@@ II-B4 Translation
To make medical LLMs widely accessible across various countries and regions, addressing language differences is crucial. Translating existing medical corpora can overcome this barrier and maximize the utilization of high-quality data. Consequently, some medical LLMs work are now focusing on ensuring the accuracy and reliability of translation results. DoctorGLM [26] prompts ChatGPT to translate the medical corpus data to the target language, and then trains a new medical translation model for learning the mapping from the source language to the target language. BiMediX [52] employs ChatGPT for English-to-Arabic translation, followed by a quality assessment using ChatGPT to ensure terminology accuracy. If translations do not meet a preset standard, an improvement process begins. This involves providing ChatGPT with the original English text, the current translation, and its evaluation score, prompting it to refine the translation for better consistency with the original. The translation quality is continuously enhanced through iterative feedback. Translations that remain unsatisfactory are reviewed and corrected by Arabic-speaking medical experts. Finally, a random sample of high-scoring translations is reviewed to ensure data quality. To maintain uniformity in language and style within the training corpus, HuatuoGPT-II [45] was inspired by back-translation to standardize the format of the training corpus by prompting ChatGPT. In contrast, the multilingual medical LLM, Apollo [54], considers cultural differences, beliefs, and taboo phrases across regions. It uses local language medical datasets without translation to avoid potential conflicts.
III Training Paradigms
@@
Transitioning from general LLMs to a medical LLMs involves three stages of continued training: continued pretraining, instruction fine-tuning, and human alignment. However, given that the general LLM has completed training and possesses some medical knowledge alongside its general knowledge, all three phases may not be necessary to train a medical LLM based on the general LLM. Tab. II presents information about the medical LLMs, including number of parameters, base model, and training stages employed. Medical LLMs learn different capabilities at various training stages, and have diverse demands on computational resources and dataset scales. We categorize the training process into four paradigms based on the combination of training stages, as depicted in Fig. 3. The characteristics and strategies of each paradigm are described in detail below.
@@
TABLE II: Detailed Information of Medical Large Language Models. Note: ”Para.” denotes parameters, ”CP” denotes continued pretraining, ”IFT” denotes instruction fine-tuning, and ”HA” denotes human alignment.
@@
Models Backbone Para. (B) CP IFT IFT Methods HA Preferred Languages Open Sources Data
@@
Figure 3: Training Paradigms. Details are provided on the training phases for each paradigm and the capabilities the model acquires at each phase.
III-A IFT
@@
Instruction fine-tuning, also termed supervised fine-tuning, refers to parameter fine-tuning of pretrained LLMs using instruction data (QA pairs). During the training process, only the loss of response content is considered. After instruction fine-tuning, medical LLMs can demonstrate strong medical instruction following and zero-shot learning ability. It can effectively grasp instructions from medical practitioners or patients, thereby enhancing doctors’ clinical efficiency and addressing patients’ needs efficiently.
@@
Medical LLMs depicted in Tab. II all employ instruction fine-tuning to enhance their capacity for addressing medical issues. This is because the general LLMs inherently possess extensive knowledge, and thus, models only require instruction fine-tuning to adapt to medical directives. Additionally, the emergence of Parameter Efficient Fine-Tuning (PEFT) accelerates the training process of instruction fine-tuning and reduces the demand for computational resources. Strengths: Instruction fine-tuning (IFT) can improve large medical language models’ performance in instruction following adherence with minimal computational resources and time consumption. Weaknesses: Without continued pretraining and human alignment, the knowledge understanding of the model may be limited and the response of the model may not accurately simulate the characteristics of a physician.
III-A1 Parameter Efficient Fine-Tuning
@@
As can be seen from Tab. II, there are two types of instruction fine-tuning methods used in medical LLMs, one is full parameter fine-tuning and the other is Parameter Efficient Fine-Tuning (PEFT) [114]. Full parameter fine-tuning involves training and updating all parameters related to the medical LLM during instruction fine-tuning. Conversely, PEFT reduces the number of model parameters requiring training, thereby reducing the performance demands of the necessary equipment for training, while ensuring that the fine-tuned model’s performance is comparable to that of full parameter fine-tuning. The most commonly used parameter-efficient fine-tuning method in the medical LLMs is Low-Rank Adaptation (LoRA) [115]. LoRA approximates parameter updates for each layer by adding low-rank decomposition matrices while freezing the original weights, thereby reducing the number of training parameters and the training time. Another commonly used method is QLoRA [116], which is a parameter quantization technique based on the LoRA method. In QLoRA, the original model parameters are quantized to 4-bit, while the newly added low-rank decomposition matrix parameters are saved as 16-bit for training. This approach further reduces the performance requirements of the training equipment and lowers the threshold for developing medical LLMs in medical institutions. In addition, there are many PEFT methods, such as Adapter Tuning [117], Prefix Tuning [118], Cpoly [119], Prompt Tuning [120] and P-tuning [121, 122], which can also reduce the performance requirements and time consumption of the training equipment. However, these methods are used less frequently than LoRA.
III-A2 Two-stage Fine-tuning Strategy
@@
Given the variety of data sources for instruction fine-tuning, each representing different aspects of medical knowledge, one-stage instruction fine-tuning may not fully utilize these diverse sources to inject medical instruction following capabilities into the model. Consequently, some efforts in developing medical LLMs have begun to adopt a two-stage instruction fine-tuning approach. DISC-MedLLM [37] adopted a two-stage training strategy for the instruction fine-tuning phase. The first stage involved training on large-scale instruction data, utilizing various datasets to inject medical domain knowledge and dialog capabilities into the model. The second stage focused on a small amount of high-quality data, generated with the assistance of ChatGPT and filtered and guided by humans, characterized by matching human behavioral preferences. The purpose of this phase is to align the behavior of the medical LLMs with human preferences. Taiyi [42] considers that the instruction dataset comprises multiple tasks, and one-stage instruction fine-tuning can lead to interference between tasks, necessitating a two-stage strategy. The first stage uses conventional NLP task datasets, such as information extraction and text classification. The second stage uses QA and dialog task datasets, training these alongside the first stage data. This two-stage training allows the model to initially focus on processing non-generative tasks and then, in the second phase, to enhance its cross-task and dialog task processing capabilities. OncoGPT [53] is first trained on Alpaca and doctor-patient communication data from online medical consultation platforms for instruction fine-tuning to gain basic medical conversational skills. It is then further trained on real doctor-patient oncology conversation data thoroughly proofread and reviewed by experts and clinicians to improve cancer information processing and conversation skills.
III-B CP-IFT
@@
Continued pretraining, also known as continued domain-adaptive pretraining, involves further pretraining of LLMs on domain-specific corpora to enrich domain knowledge [123]. The CP-IFT paradigm is that the general LLM initially performs continued pretraining using unstructured medical data to enhance its medical knowledge, adapt to the specific terminology, language style, and text structure of the medical domain, and gain a deeper understanding of medical expertise and complex concepts. It then performs instruction fine-tuning on medical instruction data to improve its ability to follow medical instructions. There exist a few studies that follow the CP-IFT paradigm to train medical LLMs on top of general LLMs. Additionally, there is a situation where some studies [44, 48, 49] perform continued pretraining of a general LLM on a medical corpus. However, since the model does not possess or has forgotten its instruction following capabilities at this time, instruction fine-tuning is conducted on the training set of the evaluation benchmark.
@@ III-B1 Model Merging
Model fusion refers to the merging of two or more models in order to improve the generalization performance of the model without additional training. Most methods focus on fusing model parameters. BioMistral merges the pretrained medical LLMs with the base model using three methods (TIES, DARE, and SLERP) for weight merging respectively, to improve the general domain capabilities of the medical LLMs. Apollo [54] uses Proxy Tuning [124] to indirectly guide the adjustment of general LLMs with a large number of parameters by utilizing the logits (outputs) of the pre-fine-tuning model (without medical knowledge injected) and the fine-tuned model (with medical knowledge injected). Experimental findings indicate that proxy tuning enhances the model’s performance on multiple language benchmarks beyond the English benchmark.
III-B2 Combining Pretraining Data with Instruction Data
@@
Incorporating pretraining data into instruction fine-tuning by converting pretraining data into the format of instruction data and jointly training it with other instruction data enhances the effectiveness and stability of training [20]. Building upon this concept, HuatuoGPT-II [45] transforms its pretrained medical data into an instruction format, creating a pre-trained instruction dataset. This dataset is then jointly trained with existing instruction data. Additionally, it proposes a priority sampling strategy to mitigate the challenges associated with mixing data sources during LLM training. Similarly, Apollo [54] rewrites the pretraining corpus into the data format of QA and adopts the same data mixing and sampling techniques as HuatuoGPT-II. This approach facilitates a smoother transition from continued pretraining to instruction fine-tuning.
III-C IFT-HA
@@
The IFT-HA paradigm adopts a two-stage approach. In the first stage, instruction fine-tuning is employed to augment the LLM’s ability to follow medical instructions and engage in medical dialogue. Subsequently, the second stage focuses on human alignment, ensuring the LLM’s responses adhere to established medical alignment principles. In general domain, human alignment aims to ensure that LLMs operate in accordance with human values, genuine intentions, and established social ethics. The main alignment principles are helpful, honest, and harmless. Within the medical domain, these core principles are further extended to encompass patient-Fridendly and Doctor-Like (professional and interactive diagnostic) [8].
@@
To achieve human alignment in the medical domain, a prevalent approach involves Reinforcement Learning from Human Feedback (RLHF) [14]. This methodology entails a two-stage process. Initially, a reward model is constructed on top of an instruction fine-tuned medical LLM, utilizing manually labeled human preference data. Subsequently, this reward model is combined with either PPO [125] or Rejection Sampling strategies [101] to perform reinforcement learning training, thereby aligning medical LLMs with predetermined principles. Considering the difficulty of collecting human preference data, HuatuoGPT, motivated by RLHF and RLAIF [126], proposes Reinforcement Learning with Mixed Feedback (RLMF). RLMF leverages ChatGPT to generate a portion of the human preference data, effectively merging the strengths of ChatGPT and real doctors. Consequently, medical LLMs are guided to align responses with both ChatGPT and medical expertise. Qilin-Med [41] adopts Direct Preference Optimization (DPO) [127], a non-reinforcement learning approach, to accomplish human alignment training for medical LLMs. Unlike reinforcement learning methods that rely on reward modeling, DPO directly establishes the connection between the decision function and the reward function, obviating the need for explicit reward modeling.
@@
Currently, human alignment training is built on models that have undergone instruction fine-tuning. The IFT-HA paradigm offers several advantages: medical LLMs exhibit enhanced instruction following capabilities, yielding more accurate responses to user queries and physician instructions. Additionally, IFT-HA fosters alignment with human values, promoting the provision of accurate and reliable medical information, clear explanations, and the avoidance of harmful suggestions while safeguarding user privacy. However, the IFT-HA paradigm presents challenges. The human alignment training process is intricate, susceptible to instability, and highly sensitive to hyperparameter selection. Furthermore, RLHF is highly dependent on the performance of the reward model, which in turn is dependent on the quality of the expensive human preference data. Consequently, developers opting for this paradigm need substantial engineering experience.
III-D CP-IFT-HA
@@
The CP-IFT-HA paradigm injects medical knowledge at different dimensions into the general LLM from three stages: continued pretraining, instruction fine-tuning and human alignment. Due to the complexity of training and large computational resources required for this paradigm, only a few medical LLM efforts have adopted this paradigm. However, the medical knowledge comprehension, medical instruction following and human preference abilities of medical LLMs are improved compared to other paradigms.
@@
Compared to the CP-IFT paradigm , the paradigm introduces the human alignment phase. This enhancement empowers medical LLMs to generate outputs that are more aligned with medical principles and human expectations, thereby reducing potentially harmful outputs and fostering greater caution when answering medical questions. In particular, the paradigm can better protect patient privacy when dealing with sensitive medical information. Compared to the IFT-HA paradigm, this CP-IFT-HA paradigm incorporates additional stages of continued pretraining, resulting in medical LLMs equipped with a more robust foundation in medical knowledge. This enhanced knowledge enables models to posses greater adaptability within the medical domain. Consequently, models using this paradigm are able to provide more accurate and insightful answers, which is especially important in medical domains that involve highly specialized and precise questions.
III-E Summary

### v2

@@ 2.2.4 Translation
To make medical LLMs widely accessible across various countries and regions, addressing language differences is crucial. Translating existing medical corpora can overcome this barrier and maximize the utilization of high-quality data. Consequently, some medical LLMs are now focusing on ensuring the accuracy and reliability of translation results. DoctorGLM [172] prompts ChatGPT to translate the medical corpus data to the target language and then trains a new medical translation model for learning the mapping from the source language to the target language. BiMediX [118] employs ChatGPT for English-to-Arabic translation, followed by a quality assessment using ChatGPT to ensure terminology accuracy. If translations do not meet a preset standard, an improvement process begins. This involves providing ChatGPT with the original English text, the current translation, and its evaluation score, prompting it to refine the translation for better consistency with the original. The translation quality is continuously enhanced through iterative feedback. Translations that remain unsatisfactory are reviewed and corrected by Arabic-speaking medical experts. Finally, a random sample of high-scoring translations is reviewed to ensure data quality. To maintain uniformity in language and style within the training corpus, HuatuoGPT-II [22] was inspired by back-translation to standardize the format of the training corpus by prompting ChatGPT. In contrast, the multilingual medical LLM, Apollo [156], considers cultural differences, beliefs, and taboo phrases across regions. It uses local language medical datasets without translation to avoid potential conflicts.
3 Training Paradigms
@@
Transferring from general to medical LLMs involves three stages of continued training: continued pretraining, instruction fine-tuning, and human alignment. However, given that the general LLM has completed training and possesses some medical knowledge alongside its general knowledge, all three stages may not be necessary to train a medical LLM based on the general LLM. Tab. 2 presents information about the medical LLMs, including the number of parameters, base model, and training stages employed. Medical LLMs learn different capabilities at various training stages and have diverse demands on computational resources and dataset scales. We categorized four paradigms according to the combination of training stages employed by existing medical LLMs, as depicted in Fig. 4. The characteristics and strategies of each paradigm are described in detail below.
@@
Table 2: Detailed Information of Medical Large Language Models. Note: "Para." denotes parameters, "CP" denotes continued pretraining, "IFT" denotes instruction fine-tuning, and "HA" denotes human alignment.
@@
Models Backbone Para. (B) CP IFT IFT Methods HA Preferred Languages Open Sources Date
@@
Hippocrates [3] LLaMA2 [146],
Mistral [70]
7 $\checkmark$ $\checkmark$ LoRA $\checkmark$ EN, CN 04/2024
MING-MOE [94] Qwen1.5-Chat [8] 1.8, 4, 7, 14 $\checkmark$ LoRA EN, CN 04/2024
Lingdan [67] Baichuan2 [176] 13 $\checkmark$ $\checkmark$ QLoRA EN, CN 04/2024
Aloe [51] Mistral [70],
LLaMA3 [5]
7, 8 $\checkmark$ Full Para. $\checkmark$ EN 05/2024
PediatricsGPT [177] Baichuan2 [176] 7,13 $\checkmark$ $\checkmark$ Full Para., LoRA $\checkmark$ CN 06/2024
Aqulia-Med [198] Aquila [170] 7 $\checkmark$ $\checkmark$ Full Para. $\checkmark$ EN, CN 06/2024
Figure 4: Training Paradigms. The training stage, the achieved capabilities, the required computing resources and the training complexity are provided for each paradigm.
3.1 IFT Paradigm
@@
Instruction fine-tuning (IFT), also termed supervised fine-tuning (SFT), refers to the parameter fine-tuning of pretrained LLMs using instruction data (e.g., QA pairs). During the training process, only the loss of response content is considered. After instruction fine-tuning, medical LLMs can demonstrate strong medical instruction following and zero-shot learning ability. It can effectively grasp instructions from medical practitioners or patients, thereby enhancing doctors’ clinical efficiency and addressing patients’ needs efficiently.
@@
Medical LLMs depicted in Tab. 2 all employ instruction fine-tuning to enhance their capacity for addressing medical issues. This paradigm leverages the extensive knowledge base inherent in general LLMs, requiring only instruction fine-tuning to adapt LLMs to specific medical tasks. Additionally, the emergence of Parameter Efficient Fine-Tuning (PEFT) [93] accelerates the training process of instruction fine-tuning and reduces the demand for computational resources. Strengths: Instruction fine-tuning can improve large medical language models’ performance in instruction following adherence with minimal computational resources and time consumption. Weaknesses: Without continued pretraining and human alignment, the knowledge understanding of the model may be limited, and its response may not accurately simulate the characteristics of the physician.
3.1.1 Parameter Efficient Fine-Tuning
@@
As can be seen from Tab. 2, there are two types of instruction fine-tuning methods used in medical LLMs: one is full parameter fine-tuning, and the other is PEFT [93]. Full parameter fine-tuning involves training and updating all parameters related to the medical LLM during instruction fine-tuning. Conversely, PEFT reduces the number of model parameters requiring training, thereby reducing the performance demands of the necessary equipment for training while ensuring that the fine-tuned model’s performance is comparable to that of full parameter fine-tuning. The most commonly used parameter-efficient fine-tuning method in the medical LLMs is Low-Rank Adaptation (LoRA) [66]. LoRA approximates parameter updates for each layer by adding low-rank decomposition matrices while freezing the original weights, thereby reducing the number of training parameters and the training time. Another commonly used method is QLoRA [33], which is a parameter quantization technique based on the LoRA method. In QLoRA, the original model parameters are quantized to 4-bit, while the newly added low-rank decomposition matrix parameters are saved as 16-bit for training. This approach further reduces the performance requirements of the training equipment and lowers the threshold for developing medical LLMs in medical institutions. In addition, there are many PEFT methods, such as Adapter Tuning [65], Prefix Tuning [90], Cpoly [152], Prompt Tuning [84] and, P-tuning [100, 99], which can also reduce the performance requirements and time consumption of the training equipment. However, these methods are used less frequently than LoRA.
3.1.2 Two-stage Fine-tuning Strategy
@@
Given the diverse sources of data reflecting various aspects of medical knowledge, a one-stage instruction fine-tuning approach may not fully exploit these distinct data sources. Consequently, some efforts in developing medical LLMs have begun to adopt a two-stage instruction fine-tuning approach. DISC-MedLLM [11] adopted a two-stage training strategy for the instruction fine-tuning stage. The first stage involved training on large-scale instruction data, utilizing various datasets to inject medical domain knowledge and dialogue capabilities into the model. The second stage focused on a small amount of high-quality data, generated with the assistance of ChatGPT and filtered and guided by humans, characterized by matching human behavioral preferences. This stage aims to align the behavior of the medical LLMs with human preferences. Taiyi [102] considers that the instruction dataset comprises multiple tasks, and one-stage instruction fine-tuning can lead to interference between tasks, necessitating a two-stage strategy. The first stage uses conventional NLP task datasets, such as information extraction and text classification. The second stage uses QA and dialogue task datasets, training these alongside the first stage data. This two-stage training allows the model to initially focus on processing non-generative tasks and then, in the second stage, enhance its cross-task and dialogue task processing capabilities. OncoGPT [69] is first trained on Alpaca and doctor-patient communication data from online medical consultation platforms for instruction fine-tuning to gain basic medical conversational skills. It is then further trained on real doctor-patient oncology conversation data thoroughly proofread and reviewed by experts and clinicians to improve cancer information processing and conversation skills.
3.2 CP-IFT Paradigm
@@
Continued pretraining, also known as continued domain-adaptive pretraining, involves further pretraining of LLMs on domain-specific corpora to enrich domain knowledge [76]. The CP-IFT paradigm is that the general LLM initially performs continued pretraining using unstructured medical data to enhance its medical knowledge, adapt to the specific terminology, language style, and text structure of the medical domain, and gain a deeper understanding of medical expertise and complex concepts. It then performs instruction fine-tuning on medical instruction data to improve its ability to follow medical instructions. A few studies follow the CP-IFT paradigm to train medical LLMs on top of general LLMs. Additionally, there are several studies [24, 79, 121] where their instruction fine-tuning is executed on the training set of the evaluation benchmarks for overcoming forgetfulness of instruction following ability.
@@ 3.2.1 Model Merging
Model Merging refers to the merging of two or more models in order to improve the generalization performance of the model without additional training. Most methods focus on fusing model parameters. BioMistral [79] merges the pretrained medical LLMs with the base model using three methods (TIES [175], DARE [182], and SLERP [129]) for weight merging respectively, to improve the general domain capabilities of the medical LLMs. Apollo [156] uses Proxy Tuning [98] to indirectly guide the adjustment of general LLMs with a large number of parameters by utilizing the logits (outputs) of the pre-fine-tuning model (without medical knowledge injected) and the fine-tuned model (with medical knowledge injected). Experimental findings indicate that proxy tuning enhances the model’s performance on multiple language benchmarks beyond the English benchmark.
3.2.2 Combining Pretraining Data with Instruction Data
@@
Transforming pretraining data into instruction data and jointly training it with existing instruction data can enhance the effectiveness and stability of the training process. [58]. Building upon this concept, HuatuoGPT-II [22] transforms its pretrained medical data into an instruction format, creating a pre-trained instruction dataset. This dataset is then jointly trained with existing instruction data. Additionally, it proposes a priority sampling strategy to mitigate the challenges associated with mixing data sources during LLM training. Similarly, Apollo [156] rewrites the pretraining corpus into the data format of QA and adopts the same data mixing and sampling techniques as HuatuoGPT-II. This approach facilitates a smoother transition from continued pretraining to instruction fine-tuning.
3.3 IFT-HA Paradigm
@@
The IFT-HA paradigm adopts a two-stage approach. In the first stage, instruction fine-tuning is employed to augment the LLM’s ability to follow medical instructions and engage in medical dialogue. Subsequently, the second stage focuses on human alignment, ensuring the LLM’s responses adhere to established medical alignment principles. In the general domain, human alignment aims to ensure that LLMs operate in accordance with human values, genuine intentions, and established social ethics. The primary alignment principles are helpful, honest, and harmless. Within the medical domain, these core principles are further extended to encompass patient-friendly and doctor-like (professional and interactive diagnostic) [188].
@@
To achieve human alignment in the medical domain, a prevalent approach is Reinforcement Learning from Human Feedback (RLHF) [114]. This methodology entails a two-stage process. Initially, a reward model is constructed on top of an instruction fine-tuned medical LLM, utilizing manually labeled human preference data. Subsequently, this reward model is combined with either PPO [128] or the rejection sampling strategy [146] to perform reinforcement learning training, thereby aligning medical LLMs with predetermined principles. Considering the difficulty of collecting human preference data, HuatuoGPT, motivated by RLHF and RLAIF [81], proposes Reinforcement Learning with Mixed Feedback (RLMF). RLMF leverages ChatGPT to generate a portion of the human preference data, effectively merging the strengths of ChatGPT and real doctors. Consequently, medical LLMs are guided to align responses with both ChatGPT and medical expertise. Qilin-Med [180] adopts Direct Preference Optimization (DPO) [122], a non-reinforcement learning approach, to accomplish human alignment training for medical LLMs. Unlike reinforcement learning methods that rely on reward modeling, DPO directly establishes the connection between the decision function and the reward function, obviating the need for explicit reward modeling.
@@
Currently, human alignment training is built on models that have undergone instruction fine-tuning. The IFT-HA paradigm offers several advantages: medical LLMs exhibit enhanced instruction following capabilities, yielding more accurate responses to user queries and physician instructions. Additionally, IFT-HA fosters alignment with human values, promoting the provision of accurate and reliable medical information, clear explanations, and the avoidance of harmful suggestions while safeguarding user privacy. However, the IFT-HA paradigm presents disadvantages. The human alignment training process is intricate, susceptible to instability, and highly sensitive to hyperparameter selection. Furthermore, RLHF is highly dependent on the performance of the reward model, which in turn is dependent on the quality of the expensive human preference data. Consequently, developers opting for this paradigm need substantial engineering experience.
3.4 CP-IFT-HA Paradigm
@@
The CP-IFT-HA paradigm injects medical knowledge at different dimensions into the general LLM in three stages: continued pretraining, instruction fine-tuning, and human alignment. Due to the complexity and substantial computational resources required, few medical LLM efforts have adopted this paradigm. However, medical LLMs trained using this paradigm demonstrate superior medical knowledge comprehension, instruction following, and alignment with human preferences compared to other paradigms [179].
@@
Compared to the CP-IFT paradigm, the paradigm introduces the human alignment stage. This enhancement empowers medical LLMs to generate outputs that are more aligned with medical principles and human expectations, thereby reducing potentially harmful outputs and fostering greater caution when answering medical questions. In particular, the paradigm can better protect patient privacy when dealing with sensitive medical information. Compared to the IFT-HA paradigm, CP-IFT-HA paradigm incorporates additional stages of continued pretraining, resulting in medical LLMs equipped with a more robust foundation in medical knowledge. The enhanced knowledge enables models to possess greater adaptability within the medical domain. Consequently, models using this paradigm are able to provide more accurate and insightful answers, which is especially important in medical domains that involve highly specialized and precise questions.
3.5 Summary